Two friends decide to meet in Chicago to attend a Cubs baseball game. Rob travels 174miles in the same time that Carl travels 156miles Rob's trip uses more interstate highways and he can average 6 mph more than Carl. What is Rob's average speed?

Answer by Cromlix(4381)

Carl's speed = x
Rob's speed = x + 6
Carl's time = 156 miles/x
Rob's time = 174/x + 6
As the time were the same
we can say:-
156/x = 174/x + 6
Cross multiply
174(x) = 156(x + 6)
174x = 156x + 936
174x - 156x = 936
18x = 936
x = 936/18
x = 52 mph
Carl's speed = 52 mph
Rob's speed = 58 mph
Proof:
156/52 = 3hrs
174/58 = 3hrs
Hope this helps.
:-)